OpenStudy (anonymous):

solve:(5x - 1)/2 + (-3x - 1)/4 = 9/2 can someone show my how to do this step by step

OpenStudy (anonymous):

(5x-1)/2 + (-3x-1)/4 = 9/2 get common denoominator 4 x 2 = 8 so now we have 2 into 8 goes 4 multiplied by (5x-1) and 4 into 8 goes 2 multiplied by (-3x-1) all over 8.lets write it out: [4(5x-1) + 2(-3x-1)] /8 = 9/2 now lets cross multiply to get rid of fractions 2 [4(5x-1)+2(-3x-1)] = 8 x 9 2[(20x-4)+(-6x-2)] = 72 2[20x-6x-4-2]=72 2[14x-6]=72 28x-12=72 28x=72+12 28x=84 x=84/28 x=3 need to check this answer to be sure its right! I have and it is correct!